# Biorthogonal measures, polymer partition functions, and random matrices

### v2

@@ 1 Introduction
$\mathbb{E}\left[\prod_{k=1}^{N}(1-\sigma(x_{k}))\right]=\sum_{k=0}^{N}\frac{(-1)^{k}}{k!}\int_{\mathbb{R}^{k}}\det\left(L_{N}(x_{j},x_{m})\right)_{j,m=1}^{k}\prod_{j=1}^{k}\sigma(x_{k}){\rm d}x_{k}.$ (1.6)
@@
Note that we allow $W_{N}$ to have other zeros in $\mathcal{S}_{N}$ than $a_{1},\ldots,a_{N}$, and that the zeros $a_{1},\ldots,a_{N}$ do not need to be distinct. The expression (1.8) is independent of the choice of ${\Sigma_{N}}$ and $\ell_{N}$, as long as the above assumptions are satisfied. Under these assumptions, we will show that the complex-valued measures (1.7) are biorthogonal: they can be written in the form (1.1), but they are not necessarily positive. We will use the term biorthogonal measure instead of biorthogonal ensemble to indicate that the measures are in general not probability measures. We also define, in analogy with (1.2), marginals for such measures,
@@
[…]
@@
$\begin{array}[]{rcl}W_{N}^{{\rm LUE+}}(z;\vec{a},\nu)&=&\displaystyle\frac{\prod_{j=1}^{N}(z-a_{j})}{(z-1)^{N+\nu}},\\ \\ W_{N}^{{\rm GUE+}}(z;\vec{a},\tau)&=&\displaystyle{\rm e}^{\tau z^{2}/2}\prod_{m=1}^{N}(z-a_{m})\\ \\ W_{N}^{{\rm GLUE+}}(z;\vec{a},\tau)&=&\displaystyle{\rm e}^{\tau z^{2}/2}\frac{\prod_{k=1}^{N}(z-a_{k})}{(1-z)^{N}}.\end{array}$ (1.18)
They are associated, respectively, to the biorthogonal ensembles given by the eigenvalues of an LUE (Laguerre Unitary Ensemble) matrix with external source, a GUE (Gaussian Unitary Ensemble) with external source, and a sum of an LUE matrix with external source and a GUE matrix. We anticipate that these three models describe the small temperature limits of three different polymer models: Log Gamma, O’Connell-Yor and O’Connell-Yor with boundary sources. More details about these small temperature limits are given at the end of the introduction.
@@
$\det(1-K)_{L^{2}(\gamma)}=\sum_{k=0}^{\infty}\frac{(-1)^{k}}{k!}\int_{\gamma^{k}}\det\left(K(z_{j},z_{m})\right)_{j,m=1}^{k}\prod_{j=1}^{k}{\rm d}z_{k}.$ (1.25)
@@
Let $W_{N},{\Sigma_{N}},\ell_{N}$ satisfy Assumption 1.1, let $\sigma:\mathbb{R}\to\mathbb{C}$ be bounded and $C^{1}$ with bounded derivative, and define $\sigma_{t}(x):=\sigma(x+t)$. Let ${\rm d}\mu_{N}$ and ${\rm d}\nu_{N,t}$ be defined by (1.7) and (1.33). If $\mu_{N}[\sigma_{t}]\neq 0$, we have the identity
@@
and prove that these measures characterize, respectively, the partition functions of the Log Gamma, the O’Connell-Yor, and the mixed polymer, see Corollaries 5.2, 6.1, and 7.3. Moreover, we also prove that in a certain “small temperature” regime, these biorthogonal measures degenerate to the ones described in Remark 1.3 and associated to matrix models, see Propositions 5.4, 6.3 and 7.6. For the Log Gamma and O’Connell-Yor polymers, this is not surprising, since the small temperature degeneration of the Log Gamma polymer is a last passage percolation model well-known to be characterized in terms of the LUE with external source [44, 20, 30], while the small temperature degeneration of the O’Connell-Yor polymer is a model of Brownian queues well-known to be characterized by the GUE with external source [8, 38]. Nevertheless, it reveals how the zero temperature limit takes place on the level of the biorthogonal measures. For the mixed polymer, to the best of our knowledge, the degeneration to the matrix model given by the sum of LUE with external source and GUE is completely new.
@@ 2 Biorthogonal structure of the measures
This proves part 3 of Theorem 1.2. In a similar way, one can obtain explicit expressions for the biorthogonal measures in cases where some, but not all, of the zeros $a_{1},\ldots,a_{N}$ are confluent.
@@ 3.2 Fredholm determinants
such that the Fredhom series $\det(1-\sigma L_{N})_{L^{2}(\mathbb{R})}$ and $\det(1-\sigma\widetilde{L}_{N})_{L^{2}(\mathbb{R})}$ are equal by definition, but the integral operator $\mathcal{L}_{N}$ defined by
$\mathcal{L}_{N}[f](x):=\int_{\mathbb{R}}\widetilde{L}_{N}(x,x^{\prime})f(x^{\prime}){\rm d}x^{\prime},$ (3.4)
@@
for sufficiently small $\epsilon>0$, with $a_{\min}$ and $a_{\max}$ as in (1.28). Then, the asymptotic behavior (2.3) for $\psi({\rm e}^{x})$ as $x\to\pm\infty$ implies indeed that $\mathcal{L}_{N}$ is a bounded linear operator on $L^{2}(\mathbb{R})$. Similarly, we define for $H_{N}^{\sigma}$ from (1.26) the conjugated kernel
@@
$\mathcal{H}_{N}^{\sigma}[f](y)=\int_{0}^{\infty}\widetilde{H}_{N}^{\sigma}(y,y^{\prime})f(y^{\prime}){\rm d}y^{\prime},$
@@
The operators $\mathcal{H}_{N}^{\sigma}$ and $\sigma\mathcal{L}_{N}$ are trace-class, and we have the identity
$\det(1-\mathcal{H}_{N}^{\sigma})_{L^{2}(0,\infty)}=\det(1-\sigma\mathcal{L}_{N})_{L^{2}(\mathbb{R})}.$ (3.7)
@@
$\mathcal{A}:L^{2}(\mathbb{R})\to L^{2}(0,\infty)\quad\text{and}\quad\mathcal{B}:L^{2}(0,\infty)\to L^{2}(\mathbb{R})$
@@
It is now straightforward to check that $\mathcal{A}$ is a Hilbert-Schmidt operator.
@@
which implies that $\mathcal{B}$ is also a Hilbert-Schmidt operator, under the condition that
@@
As compositions of Hilbert-Schmidt operators, $\mathcal{A}\mathcal{B}$ and $\mathcal{B}\mathcal{A}$ are trace-class operators, and we have
$\det(1-\mathcal{A}\mathcal{B})_{L^{2}(0,\infty)}=\det(1-\mathcal{B}\mathcal{A})_{L^{2}(\mathbb{R})}.$ (3.8)
It is easy to see that $\mathcal{H}_{N}^{\sigma}=\mathcal{A}\mathcal{B}$ by comparing the kernels of both operators, and we will now prove that $\sigma\mathcal{L}_{N}=\mathcal{B}\mathcal{A}$, which implies the result.
Using the integral representation of $\Psi_{1}$ and $\Psi_{2}$ from (1.27), we compute the kernel of the operator $\mathcal{B}\mathcal{A}$ acting on $L^{2}(\mathbb{R})$ as follows:
@@
such that indeed $\mathcal{B}\mathcal{A}=\sigma\mathcal{L}_{N}$. ∎
@@
We constructed $\mathcal{L}_{N}$ and $\mathcal{H}_{N}^{\sigma}$ in such a way that
$\det(1-\sigma\mathcal{L}_{N})_{L^{2}(0,\infty)}=\det(1-\sigma L_{N})_{L^{2}(\mathbb{R})},\qquad\det(1-\mathcal{H}_{N}^{\sigma})_{L^{2}(0,\infty)}=\det(1-H_{N}^{\sigma})_{L^{2}(\mathbb{R})},$
@@
$\det\left(1-\mathcal{H}_{N}^{\sigma_{t}}\right)_{L^{2}(\mathbb{R})}=\det\left(1+\mathcal{K}_{N,t}\right)_{L^{2}({\Sigma_{N}})},$ (3.9)
with $\mathcal{K}_{N,t}$ the integral operator acting on $L^{2}(\Sigma_{N})$ with kernel $K_{N,t}$ defined in (1.29).
@@
Using the fact that $\frac{1}{v-u^{\prime}}=\displaystyle\int_{0}^{\infty}{\rm d}y\,{\rm e}^{-y(v-u^{\prime})}$ when $\mathrm{Re}(v-u^{\prime})>0$, we can write $K_{N,t}$ from (1.29) as the composition of two Hilbert-Schmidt operators, $\mathcal{K}_{N,t}=\mathcal{C}\mathcal{D}$, with
$\mathcal{C}:L^{2}(0,\infty)\to L^{2}({\Sigma_{N}})\;\text{and}\;\mathcal{D}:L^{2}({\Sigma_{N}})\to L^{2}(0,\infty)$
@@
where $\ell_{N}$ is the vertical line with real part $c$, with $a_{\max}<c<a_{\min}+1$, and where $\Sigma_{N}$ lies at the left of $\ell_{N}$. It is then straightforward to verify that $\mathcal{D}$ is a Hilbert-Schmidt operator. For $\mathcal{C}$, the same is true, but note that we need here that $10<\mathrm{Re}\,(v-u)<$ 1 for every $v\in\ell_{N}$, $u\in{\Sigma_{N}}$ in order to avoid zeros of $\sin\pi(u-v)$.
Now we use the fact that $\det(1+\mathcal{C}\mathcal{D})_{L^{2}({\Sigma_{N}})}=\det(1+\mathcal{D}\mathcal{C})_{L^{2}(0,\infty)}$, and compute the kernel of $\mathcal{D}\mathcal{C}$ as (note that the minus sign in front compensates the modified sign of the argument of $\sin$)
@@ 4 Logarithmic derivatives and deformed biorthogonal measures
In this section, we will prove Theorem 1.10. The proof of this result is essentially a special case of a more general result in [26]. Besides Assumption 1.1, we assume here that $\sigma:\mathbb{R}\to\mathbb{C}$ is $C^{1}$, bounded, and with bounded derivative, and that $t$ is such that $\mu_{N}[\sigma_{t}]=\det(1-\sigma_{t}\mathcal{L}_{N})_{L^{2}(\mathbb{R})}\neq 0$, with $\sigma_{t}(x)=\sigma(x+t)$.
@@
with $F_{k}(x)=h(x)\psi_{k}({\rm e}^{x})$ and $G_{k}(x)={\rm e}^{a_{k}x}/h(x)$. The asymptotic behavior (2.3) of $\psi_{k}$ and the definition (3.5) imply that $F_{k}(x)$ and $G_{k}(x)$ decay rapidly as $x\to\pm\infty$. Consequently, the operator $\mathcal{L}_{N}$ is a bounded linear operator, which is moreover finite rank and thus trace-class. By the Jacobi identity, we have
$\frac{{\rm d}}{{\rm d}t}\log\mu_{N}[\sigma_{t}]=\frac{{\rm d}}{{\rm d}t}\log\det(1-\sigma_{t}\mathcal{L}_{N})_{L^{2}(\mathbb{R})}=-\mathrm{Tr}\,\left(\left(\frac{{\rm d}}{{\rm d}t}\sigma_{t}\right)\mathcal{L}_{N}(1-\sigma_{t}\mathcal{L}_{N})^{-1}\right)_{L^{2}(\mathbb{R})},$ (4.2)
where we observe that $(\frac{{\rm d}}{{\rm d}t}\sigma_{t})\mathcal{L}_{N}(1-\sigma_{t}\mathcal{L}_{N})^{-1}=\sigma_{t}^{\prime}\mathcal{L}_{N}(1-\sigma_{t}\mathcal{L}_{N})^{-1}$ is a trace-class operator if $\sigma_{t}^{\prime}$ is bounded. We thus have
$\frac{{\rm d}}{{\rm d}t}\log\mu_{N}[\sigma_{t}]=-\mathrm{Tr}\,\left(\sigma_{t}^{\prime}\mathcal{L}_{N}(1-\sigma_{t}\mathcal{L}_{N})^{-1}\right)_{L^{2}(\mathbb{R})}.$ (4.3)
By definition (3.4) and by Proposition 2.2, we directly verify that $\mathcal{L}_{N}^{2}=\mathcal{L}_{N}$, such that $\mathcal{L}_{N}$ is a rank $N$ (not necessarily Hermitian) projection operator. We denote $\mathcal{F}_{N}$ for the linear span of $F_{1},\ldots,F_{N}$ and $\mathcal{G}_{N}$ for the linear span of $G_{1},\ldots,G_{N}$.
@@
$\mathcal{L}_{N}$ is the unique projection operator on $L^{2}(\mathbb{R})$ with range $\mathcal{F}_{N}$ and kernel $\mathcal{G}_{N}^{\perp}$.
@@
It is straightforward to verify from (3.4)–(3.3) and (4.1) that $\mathcal{L}_{N}[f]=f$ for $f\in\mathcal{F}_{N}$ and that $\mathcal{L}_{N}[f]=0$ for $f\in\mathcal{G}_{N}^{\perp}$. Moreover, $\mathcal{F}_{N}\oplus\mathcal{G}_{N}^{\perp}=L^{2}(\mathbb{R})$. Indeed, suppose that this is not the case; then there exists a non-zero $f\in L^{2}(\mathbb{R})$ such that $f\perp h$ for every $h\in\mathcal{F}_{N}\oplus\mathcal{G}_{N}^{\perp}$. Consequently, $f\in\mathcal{F}_{N}^{\perp}\cap\mathcal{G}_{N}$. but $\mathcal{F}_{N}^{\perp}\cap\mathcal{G}_{N}=\{0\}$, by the orthogonality relation $\int_{\mathbb{R}}F_{j}(x)G_{k}(x){\rm d}x=\delta_{jk}$, which gives a contradiction. Using the above properties, we obtain that $\mathcal{L}_{N}$ is the unique projection with range $\mathcal{F}_{N}$ and kernel $\mathcal{G}_{N}^{\perp}$. ∎
Let us now consider a kernel $M_{t}(x,x^{\prime})$ of the operator $\mathcal{L}_{N}(1-\sigma_{t}\mathcal{L}_{N})^{-1}$ acting on $L^{2}(\mathbb{R})$, such that
$\mathcal{L}_{N}(1-\sigma_{t}\mathcal{L}_{N})^{-1}[f](x)=\int_{\mathbb{R}}M_{t}(x,x^{\prime})f(x^{\prime}){\rm d}x^{\prime}.$
@@
$\mathcal{P}_{t}[g](x)=\mathcal{L}_{N}(1-\sigma_{t}\mathcal{L}_{N})^{-1}[(1-\sigma_{t})g]=\int_{\mathbb{R}}M_{t}(x,x^{\prime})g(x^{\prime})(1-\sigma_{t}(x^{\prime})){\rm d}x^{\prime}.$ (4.4)
@@
Let $g\in\mathcal{F}_{N}$, such that $(1-\sigma_{t})g=(1-\sigma_{t}\mathcal{L}_{N})g$. Then,
$\mathcal{P}_{t}[g]=\mathcal{L}_{N}(1-\sigma_{t}\mathcal{L}_{N})^{-1}[(1-\sigma_{t})g]=\mathcal{L}_{N}[g]=g.$
Finally, let $g\in\mathcal{G}_{N}^{\perp}$, such that $\mathcal{L}_{N}[(1-\sigma_{t})g]=0$ and $(1-\sigma_{t})g=\left(1-\sigma_{t}\mathcal{L}_{N}\right)[(1-\sigma_{t})g]$. Thus,
$\mathcal{P}_{t}[g]=\mathcal{L}_{N}[(1-\sigma_{t})g]=0.$
@@
$\mathcal{P}_{t}[g]=\mathcal{L}_{N}(1-\sigma_{t}\mathcal{L}_{N})^{-1}[(1-\sigma_{t})g]=\int_{\mathbb{R}}M_{t}(x,x^{\prime})g(x^{\prime})(1-\sigma_{t}(x^{\prime})){\rm d}x,$
@@
where $\widetilde{F}_{1},\ldots,\widetilde{F}_{N}$ span the range $\mathcal{F}_{N}$ of $\mathcal{L}_{N}$, and $\widetilde{G}_{1},\ldots,\widetilde{G}_{N}$ span $\mathcal{G}_{N}$, since $\mathcal{G}_{N}^{\perp}$ (in $L^{2}(\mathbb{R},(1-\sigma_{t}(x)){\rm d}x)$) is the kernel of $\mathcal{P}_{t}$. The biorthogonal measure
@@
$\frac{{\rm d}}{{\rm d}t}\log\mu_{N}[\sigma_{t}]=-\mathrm{Tr}\,\left(\sigma_{t}^{\prime}\mathcal{L}_{N}(1-\sigma_{t}\mathcal{L}_{N})^{-1}\right)_{L^{2}(\mathbb{R})}=\int_{\mathbb{R}}\sigma_{t}^{\prime}(x)M_{t}(x,x){\rm d}x,$

### v3

@@ 1 Introduction
$\mathbb{E}\left[\prod_{k=1}^{N}(1-\sigma(x_{k}))\right]=\sum_{k=0}^{N}\frac{(-1)^{k}}{k!}\int_{\mathbb{R}^{k}}\det\left(L_{N}(x_{j},x_{m})\right)_{j,m=1}^{k}\prod_{j=1}^{k}\sigma(x_{j}){\rm d}x_{j}.$ (1.6)
@@
Note that we allow $W_{N}$ to have other zeros in $\mathcal{S}_{N}$ than $a_{1},\ldots,a_{N}$, and that the zeros $a_{1},\ldots,a_{N}$ do not need to be distinct. The expression (1.8) is independent of the choice of ${\Sigma_{N}}$ and $\ell_{N}$, as long as the above assumptions are satisfied. Under these assumptions, we will show that the complex-valued measures (1.7) are biorthogonal: they can be written in the form (1.1). We will use the term biorthogonal measure instead of biorthogonal ensemble to indicate that the measures are in general not probability measures. All the examples we study, nevertheless, are real-valued, even if not necessarily positive. We also define, in analogy with (1.2), marginals for such measures,
@@
    $L_{N}(x,x^{\prime})=\sum_{m=1}^{N}{\rm e}^{a_{m}x^{\prime}}\psi_{m}({\rm e}^{x}),\quad\text{where}\quad\psi_{m}(y):=\frac{1}{2\pi\mathrm{i}W_{N}^{\prime}(a_{m})}\int_{\ell_{N}}\frac{W_{N}(v)y^{-v}}{v-a_{m}}{\rm d}v,$ (1.12)
@@
$\begin{array}[]{rcl}W_{N}^{{\rm LUE+}}(z;{\vec{b}},\nu)&=&\displaystyle\frac{\prod_{j=1}^{N}(z-{b_{j}})}{(z-1)^{N+\nu}},\\ \\ W_{N}^{{\rm GUE+}}(z;\vec{a},\tau)&=&\displaystyle{\rm e}^{\tau z^{2}/2}\prod_{j=1}^{N}(z-a_{j})\\ \\ W_{N}^{{\rm GLUE+}}(z;{\vec{b}},\tau)&=&\displaystyle{\frac{z^{N}{\rm e}^{\tau z^{2}/2}}{\prod_{j=1}^{N}(z-1+b_{j})}}.\end{array}$ (1.18)
They are associated, respectively, to the biorthogonal ensembles given by the eigenvalues of a LUE (Laguerre Unitary Ensemble) matrix with external source, a GUE (Gaussian Unitary Ensemble) with external source, and a sum of a LUE matrix with external source and a GUE matrix. As such, the choices of $W$ given above define positive-valued measures. We anticipate that these three models describe the small temperature limits of three different polymer models: Log Gamma, O’Connell-Yor and O’Connell-Yor with boundary sources. More details about these small temperature limits are given at the end of the introduction.
@@
$\det(1-K)_{L^{2}(\gamma)}=\sum_{k=0}^{\infty}\frac{(-1)^{k}}{k!}\int_{\gamma^{k}}\det\left(K(z_{j},z_{m})\right)_{j,m=1}^{k}\prod_{j=1}^{k}{\rm d}z_{j}.$ (1.25)
@@
Let $W_{N},{\Sigma_{N}},\ell_{N}$ satisfy Assumption 1.1, let $\sigma:\mathbb{R}\to\mathbb{C}$ be bounded and $C^{1}$ with bounded derivative, and define $\sigma_{t}(x):=\sigma(x+t)$. Let ${\rm d}\mu_{N}$ and ${\rm d}\nu_{N,t}$ be defined by (1.7) and (1.33), and $\mu_{N}[\sigma]$ defined by (1.24). If $\mu_{N}[\sigma_{t}]\neq 0$, we have the identity
@@
and prove that the associated measures are real-valued (but not necessarily positive) and characterize, respectively, the partition functions of the Log Gamma, the O’Connell-Yor, and the mixed polymer, see Corollaries 5.2, 6.1, and 7.3. Moreover, we also prove that in a certain “small temperature” regime, these biorthogonal measures degenerate to the ones described in Remark 1.3 and associated to matrix models, see Propositions 5.5, 6.4 and 7.7. For the Log Gamma and O’Connell-Yor polymers, this is not surprising, since the small temperature degeneration of the Log Gamma polymer is a last passage percolation model well-known to be characterized in terms of the LUE with external source [44, 20, 30], while the small temperature degeneration of the O’Connell-Yor polymer is a model of Brownian queues well-known to be characterized by the GUE with external source [8, 38]. Nevertheless, it reveals how the zero temperature limit takes place on the level of the biorthogonal measures. For the mixed polymer, to the best of our knowledge, the degeneration to the matrix model given by the sum of LUE with external source and GUE is completely new.
@@ 2 Biorthogonal structure of the measures
This proves part 3 of Theorem 1.2.
In a similar way, one can obtain explicit expressions for the biorthogonal measures in cases where some, but not all, of the zeros $a_{1},\ldots,a_{N}$ are confluent.
@@ 3.2 Fredholm determinants
such that the Fredhom series $\det(1-\sigma L_{N})_{L^{2}(\mathbb{R})}$ and $\det(1-\sigma\widetilde{L}_{N})_{L^{2}(\mathbb{R})}$ are equal by definition, but the integral operator $\widetilde{L}_{N}$ defined by
$\widetilde{L}_{N}[f](x):=\int_{\mathbb{R}}\widetilde{L}_{N}(x,x^{\prime})f(x^{\prime}){\rm d}x^{\prime},$ (3.4)
@@
for sufficiently small $\epsilon>0$, with $a_{\min}$ and $a_{\max}$ as in (1.28). Then, the asymptotic behavior (2.3) for $\psi({\rm e}^{x})$ as $x\to\pm\infty$ implies indeed that $\widetilde{L}_{N}$ is a bounded linear operator on $L^{2}(\mathbb{R})$. Similarly, we define for $H_{N}^{\sigma}$ from (1.26) the conjugated kernel
@@
${\widetilde{H}_{N}^{\sigma}}[f](y)=\int_{0}^{\infty}\widetilde{H}_{N}^{\sigma}(y,y^{\prime})f(y^{\prime}){\rm d}y^{\prime},$
@@
The operators ${\widetilde{H}_{N}^{\sigma}}$ and $\sigma\widetilde{L}_{N}$ are trace-class, and we have the identity
$\det(1-{\widetilde{H}_{N}^{\sigma}})_{L^{2}(0,\infty)}=\det(1-\sigma\widetilde{L}_{N})_{L^{2}(\mathbb{R})}.$ (3.7)
@@
${A:L^{2}(\mathbb{R})\to L^{2}(0,\infty)\quad\text{and}\quad B:L^{2}(0,\infty)\to L^{2}(\mathbb{R})}$
@@
It is now straightforward to check that ${A}$ is a Hilbert-Schmidt operator.
@@
which implies that ${B}$ is also a Hilbert-Schmidt operator, under the condition that
@@
As compositions of Hilbert-Schmidt operators, $AB$ and $BA$ are trace-class operators, and we have
$\det(1-AB)_{L^{2}(0,\infty)}=\det(1-BA)_{L^{2}(\mathbb{R})}.$ (3.8)
It is easy to see that $\widetilde{H}_{N}^{\sigma}=AB$ by comparing the kernels of both operators, and we will now prove that $\sigma\widetilde{L}_{N}=BA$, which implies the result.
Using the integral representation of $\Psi_{1}$ and $\Psi_{2}$ from (1.27), we compute the kernel of the operator ${BA}$ acting on $L^{2}(\mathbb{R})$ as follows:
@@
such that indeed $BA=\sigma\widetilde{L}_{N}$. ∎
@@
We constructed the Hilbert-Schmidt operators $\widetilde{L}_{N}$ and $\widetilde{H}_{N}^{\sigma}$ in such a way that
$\det(1-\sigma\widetilde{L}_{N})_{L^{2}(\mathbb{R})}=\det(1-\sigma L_{N})_{L^{2}(\mathbb{R})},\qquad\det(1-\widetilde{H}_{N}^{\sigma})_{L^{2}(0,\infty)}=\det(1-H_{N}^{\sigma})_{L^{2}(0,\infty)},$
@@
${\det\left(1-\widetilde{H}_{N}^{\sigma_{t}}\right)_{L^{2}(\mathbb{R})}=\det\left(1+K_{N,t}\right)_{L^{2}({\Sigma_{N}})}},$ (3.9)
with $K_{N,t}$ the integral operator acting on $L^{2}(\Sigma_{N})$ with kernel $K_{N,t}$ defined in (1.29).
@@
Using the fact that $\frac{1}{v-u^{\prime}}=\displaystyle\int_{0}^{\infty}{\rm d}y\,{\rm e}^{-y(v-u^{\prime})}$ when $\mathrm{Re}(v-u^{\prime})>0$, we can write $K_{N,t}$ from (1.29) as the composition of two Hilbert-Schmidt operators, $K_{N,t}=CD$, with
${C:L^{2}(0,\infty)\to L^{2}({\Sigma_{N}})\;\text{and}\;D:L^{2}({\Sigma_{N}})\to L^{2}(0,\infty)}$
@@
where $\ell_{N}$ is the vertical line with real part $c$, with $a_{\max}<c<a_{\min}+1$, and where $\Sigma_{N}$ lies at the left of $\ell_{N}$. It is then straightforward to verify that ${D}$ is a Hilbert-Schmidt operator. For ${C}$, the same is true, but note that we need here that $10<\mathrm{Re}\,(v-u)<$ 1 for every $v\in\ell_{N}$, $u\in{\Sigma_{N}}$ in order to avoid zeros of $\sin\pi(u-v)$.
Now we use the fact that $\det(1+CD)_{L^{2}({\Sigma_{N}})}=\det(1+DC)_{L^{2}(0,\infty)}$, and compute the kernel of $DC$ as (note that the minus sign in front compensates the modified sign of the argument of $\sin$)
@@ 4 Logarithmic derivatives and deformed biorthogonal measures
In this section, we will prove Theorem 1.10. The proof of this result is essentially a special case of a more general result in [26]. Besides Assumption 1.1, we assume here that $\sigma:\mathbb{R}\to\mathbb{C}$ is $C^{1}$, bounded, and with bounded derivative, and that $t$ is such that
$\mu_{N}[\sigma_{t}]=\det(1-\sigma_{t}L_{N})_{L^{2}(\mathbb{R})}=\det(1-\sigma_{t}\widetilde{L}_{N})_{L^{2}(\mathbb{R})}\neq 0,$
with $\sigma_{t}(x)=\sigma(x+t)$.
@@
with $F_{k}(x)=h(x)\psi_{k}({\rm e}^{x})$ and $G_{k}(x)={\rm e}^{a_{k}x}/h(x)$. The asymptotic behavior (2.3) of $\psi_{k}$ and the definition (3.5) imply that $F_{k}(x)$ and $G_{k}(x)$ decay rapidly as $x\to\pm\infty$. Consequently, the operator $\widetilde{L}_{N}$ is a bounded linear operator, which is moreover finite rank and thus trace-class. By the Jacobi identity, we have
$\frac{{\rm d}}{{\rm d}t}\log\mu_{N}[\sigma_{t}]=\frac{{\rm d}}{{\rm d}t}\log\det(1-\sigma_{t}\widetilde{L}_{N})_{L^{2}(\mathbb{R})}=-\mathrm{Tr}\,\left(\left(\frac{{\rm d}}{{\rm d}t}\sigma_{t}\right)\widetilde{L}_{N}(1-\sigma_{t}\widetilde{L}_{N})^{-1}\right)_{L^{2}(\mathbb{R})},$ (4.2)
where we observe that $(\frac{{\rm d}}{{\rm d}t}\sigma_{t})\widetilde{L}_{N}(1-\sigma_{t}\widetilde{L}_{N})^{-1}=\sigma_{t}^{\prime}\widetilde{L}_{N}(1-\sigma_{t}\widetilde{L}_{N})^{-1}$ is a trace-class operator if $\sigma_{t}^{\prime}$ is bounded. We thus have
$\frac{{\rm d}}{{\rm d}t}\log\mu_{N}[\sigma_{t}]=-\mathrm{Tr}\,\left(\sigma_{t}^{\prime}\widetilde{L}_{N}(1-\sigma_{t}\widetilde{L}_{N})^{-1}\right)_{L^{2}(\mathbb{R})}.$ (4.3)
By definition (3.4) and by Proposition 2.2, we directly verify that $\widetilde{L}_{N}^{2}=\widetilde{L}_{N}$, such that $\widetilde{L}_{N}$ is a rank $N$ (not necessarily Hermitian) projection operator. We denote $\mathcal{F}_{N}$ for the linear span of $F_{1},\ldots,F_{N}$ and $\mathcal{G}_{N}$ for the linear span of $G_{1},\ldots,G_{N}$.
@@
$\widetilde{L}_{N}$ is the unique projection operator on $L^{2}(\mathbb{R})$ with range $\mathcal{F}_{N}$ and kernel $\mathcal{G}_{N}^{\perp}$.
@@
It is straightforward to verify from (3.3)–(3.4) and (4.1) that $\widetilde{L}_{N}[f]=f$ for $f\in\mathcal{F}_{N}$ and that $\widetilde{L}_{N}[f]=0$ for $f\in\mathcal{G}_{N}^{\perp}$. Moreover, $\mathcal{F}_{N}\oplus\mathcal{G}_{N}^{\perp}=L^{2}(\mathbb{R})$. Indeed, suppose that this is not the case; then there exists a non-zero $f\in L^{2}(\mathbb{R})$ such that $f\perp h$ for every $h\in\mathcal{F}_{N}\oplus\mathcal{G}_{N}^{\perp}$. Consequently, $f\in\mathcal{F}_{N}^{\perp}\cap\mathcal{G}_{N}$. but $\mathcal{F}_{N}^{\perp}\cap\mathcal{G}_{N}=\{0\}$, by the orthogonality relation $\int_{\mathbb{R}}F_{j}(x)G_{k}(x){\rm d}x=\delta_{jk}$, which gives a contradiction. Using the above properties, we obtain that $\widetilde{L}_{N}$ is the unique projection with range $\mathcal{F}_{N}$ and kernel $\mathcal{G}_{N}^{\perp}$. ∎
Let us now consider a kernel $M_{t}(x,x^{\prime})$ of the operator $\widetilde{L}_{N}(1-\sigma_{t}\widetilde{L}_{N})^{-1}$ acting on $L^{2}(\mathbb{R})$, such that
$\widetilde{L}_{N}(1-\sigma_{t}\widetilde{L}_{N})^{-1}[f](x)=\int_{\mathbb{R}}M_{t}(x,x^{\prime})f(x^{\prime}){\rm d}x^{\prime}.$
@@
$\mathcal{P}_{t}[g](x)=\widetilde{L}_{N}(1-\sigma_{t}\widetilde{L}_{N})^{-1}[(1-\sigma_{t})g]=\int_{\mathbb{R}}M_{t}(x,x^{\prime})g(x^{\prime})(1-\sigma_{t}(x^{\prime})){\rm d}x^{\prime}.$ (4.4)
@@
Let $g\in\mathcal{F}_{N}$, such that $(1-\sigma_{t})g=(1-\sigma_{t}\widetilde{L}_{N})g$. Then,
$\mathcal{P}_{t}[g]=\widetilde{L}_{N}(1-\sigma_{t}\widetilde{L}_{N})^{-1}[(1-\sigma_{t})g]=\widetilde{L}_{N}[g]=g.$
Finally, let $g\in\mathcal{G}_{N}^{\perp}$, such that $\widetilde{L}_{N}[(1-\sigma_{t})g]=0$ and $(1-\sigma_{t})g=\left(1-\sigma_{t}\widetilde{L}_{N}\right)[(1-\sigma_{t})g]$. Thus,
$\mathcal{P}_{t}[g]=\widetilde{L}_{N}[(1-\sigma_{t})g]=0.$
@@
$\mathcal{P}_{t}[g]=\widetilde{L}_{N}(1-\sigma_{t}\widetilde{L}_{N})^{-1}[(1-\sigma_{t})g]=\int_{\mathbb{R}}M_{t}(x,x^{\prime})g(x^{\prime})(1-\sigma_{t}(x^{\prime})){\rm d}x,$
@@
where $\widetilde{F}_{1},\ldots,\widetilde{F}_{N}$ span the range $\mathcal{F}_{N}$ of $\widetilde{L}_{N}$, and $\widetilde{G}_{1},\ldots,\widetilde{G}_{N}$ span $\mathcal{G}_{N}$, since $\mathcal{G}_{N}^{\perp}$ (in $L^{2}(\mathbb{R},(1-\sigma_{t}(x)){\rm d}x)$) is the kernel of $\mathcal{P}_{t}$. The biorthogonal measure
@@
$\frac{{\rm d}}{{\rm d}t}\log\mu_{N}[\sigma_{t}]=-\mathrm{Tr}\,\left(\sigma_{t}^{\prime}\widetilde{L}_{N}(1-\sigma_{t}\widetilde{L}_{N})^{-1}\right)_{L^{2}(\mathbb{R})}=\int_{\mathbb{R}}\sigma_{t}^{\prime}(x)M_{t}(x,x){\rm d}x,$